OpenStudy (anonymous):

find the general solution of Ux = 0, where U := U(x,y), and Explain why when it is subject to the initial condition U(x,0) = Uo(x) is not well-posed

OpenStudy (anonymous):

@Zarkon can you help me please with this question?

OpenStudy (anonymous):

I try to solve it but I got stuck; here is what i did: I used the method of characteristic ; dx/ds=1 so x=s+c and we usually use the initial condition x(0)=x' so c=x' and dy/ds=0 so y=0 and by the initial condition y(0)=0 so c=0 so y=0, now u (x,y)= u (x,0) = uo (x')= uo (x - s), how to find s here in order to get the solution u (x,y) ???

OpenStudy (kainui):

Does Ux=0 mean \(U*x = 0\) or \(\frac{\partial U}{\partial x} = 0\)?

OpenStudy (kainui):

I guess I'm just confused since the second case looks like how you'd see a problem you'd attack with method of characteristics but uhh, It just seems like it would be unnecessary to solve that way since you can immediately integrate it. \[\frac{\partial U}{\partial x} = 0\] Since it's 0 it must be a function of y, since it doesn't depend on x. Normally this would be a constant if it was single variable though. \[U(x,y) = f(y)\]

OpenStudy (kainui):

Yeah this must be it, because this makes sense why they'd say \[U(x,0)=U_0(x)\] is not well posed since \[U_0(x)=f(y)\] So the initial conditions in this sense would mean that \(f(y)=C\) must be a constant since x and y are independent variables but equal everywhere and some condition like that would be meaningless haha.

OpenStudy (anonymous):

if it is not well posed means there is either no solution or the solution is not unique!

OpenStudy (anonymous):

i don't know i'm still a little confuse!

OpenStudy (anonymous):

why is it not well posed then? it is because the solutions depend on y so it is not unique?

OpenStudy (kainui):

Well by solving the differential equation we have: \[U(x,y)=f(y)\] So now imagine I give you some condition that \(U(x,0)=x^3\) as a specific case. Plugging in gives us: \[x^3=f(0)\] But f(0) is a constant and doesn't depend on x! So we have a contradiction.

OpenStudy (kainui):

So to specifically answer your question: "if it is not well posed means there is either no solution or the solution is not unique!" This means given an arbitrary \(U(x,0)\) there is no solution.

OpenStudy (kainui):

There's only a special case of when this works, and that's only when the initial function of x is equal to a constant \(U(x,0)=C\) (so really it's not even a function of x hah)

OpenStudy (anonymous):

thanks